OpenStudy (anonymous):

Find all solutions to the equation. 7 sin^2x - 14 sin x + 2 = -5

OpenStudy (anonymous):

I have no clue on what to do....

OpenStudy (anonymous):

Well if we make sin(x) =y then we can simplify this statement

OpenStudy (anonymous):

\[\Large 7(y^2)-14(y)+2=-5\]

OpenStudy (anonymous):

Okay, question. Why did you make sin(x)=y?

OpenStudy (anonymous):

If you move the -5 you now have it set to 0 \[\Large 7(y^2)-14(y)+7=0\]

OpenStudy (anonymous):

Oh i made it equal yo y because its much simpler to work with, just as long as we put it back at the end, there's no rules saying you can't do it, its really substitituion

OpenStudy (anonymous):

\[7\sin ^{2}(x) - 14\sin(x) + 2 = -5\] \[7\sin ^{2}(x) - 14\sin(x)= -7\] \[-\sin ^{2}(x) + 2\sin(x) = 1\] \[-\sin ^{2}(x) + 2\sin(x) - 1 = 0\] \[(-\sin(x) + 1)(\sin(x) - 1)\]

OpenStudy (anonymous):

Oh i made it equal to y because its much simpler to work with, just as long as we put it back at the end, there's no rules saying you can't do it, its really substitituion

OpenStudy (anonymous):

How did you get -sin^2x+2sinx=1?

OpenStudy (anonymous):

Divide by -7

OpenStudy (anonymous):

If you can't visually do what he's doing, he's treating sin(x) like it was y by factoring it like that but some ppl can't do that so they can use the sub. method, anyways Im going to continue my explanation

OpenStudy (anonymous):

OH BY 7

OpenStudy (anonymous):

Yup

OpenStudy (anonymous):

I'm confused on how you factored it

OpenStudy (anonymous):

oh just kidding

OpenStudy (anonymous):

Dividing both sides by 7 you get now \[\Large y^2-2y+1=0~~~~~~~~(y-1)^2=0\]

OpenStudy (anonymous):

I understand it now! Thank you all(:

OpenStudy (anonymous):

You're welcome

OpenStudy (anonymous):

sqrt both sides: y-1=0 replacing y with sin(x) again sin(x)-1=0 sin(x)=0 YW

OpenStudy (anonymous):

-sin(x) = -1 sin(x) = 1 sin(x) = 1 -2pi <= x <= 2pi x = pi/2

OpenStudy (anonymous):

sqrt both sides: y-1=0 replacing y with sin(x) again sin(x)-1=0 sin(x)=1 And it asked for all solutions so it wouldn't be in the interval -2pi<x<2pi it woul be x= pi/2+2pi(n) where n is any intger YW

OpenStudy (anonymous):

Well yea, if you included co terminal angles